. A locus containing three alleles A, B and C are found in a population known to be in Hardy Weinberg Equilibrium. What is the combined frequency of all the heterozygotes within this population if the frequency of the A allele is 0.6 and the B allele is 0.1?

a) 0.14

b) 0.60

c) 0.12

d) 0.54

e) Not enough information
Answer #1

According to Hardy-Weinberg equilibrium, p2 + 2pq + q2 =1, where p and q are the frequency of two alleles. and 2pq is the frequency of the heterozygotes.

As the frequency of two alleles is 0.6 and 0.1. So, in this way, the frequency of heterozygote is 2pq= 2*0.6*0.1 = 0.12.

So, the frequency of heterozygotes is 0.12

Therefore, the correct option is (c).

Hope it helps, Good luck !!!!

PLEASE DO PRESS LIKE :)